A stone is thrown horizontally at 20 m/s from the top of a cliff 63 m high. How fast is it moving the instant before it hits the ground?


1
Expert's answer
2022-02-18T15:53:52-0500

Given:

"v_0=20\:\rm m/s"

"H=63\:\rm m"


The law of conservation of energy says

"\frac{mv_0^2}{2}+mgH=\frac{mv^2}{2}"

"v=\sqrt{v_0^2+2gH}"

"v=\sqrt{20^2+2*9.8*63}=40.4\:\rm m/s"